OpenStudy (anonymous):

3(x-6)+8=5(x+2)-8 so i am having a few problems with this equation

OpenStudy (anonymous):

i got x=4

OpenStudy (anonymous):

First distribute: 3(x-6)+8=5(x+2)-8 3x-18+8=5x+10-8 Combine terms 3x-10=5x+2 -12=2x -6=x

OpenStudy (anonymous):

so what would the rule be onconverting the numbers be

OpenStudy (anonymous):

thank you by the way

OpenStudy (anonymous):

You're welcome, but I'm not sure what you are asking about converting the numbers....

OpenStudy (anonymous):

negative to positive

OpenStudy (anonymous):

If you divide a negative by a positive (like -12/2) you get a negative (-6)

OpenStudy (anonymous):

3x-10=5x+2 what about the -10 and the positive 2

OpenStudy (anonymous):

I subtracted 2 from both sides 3x-10=5x+2 3x-12=5x and then subtracted 3x from both sides -12=2x I should have made that more clear in my first post, sorry.

OpenStudy (anonymous):

i thank you very much for your help

OpenStudy (anonymous):

You're very welcome. There is one type of person I truly enjoy helping and that's the ones that ask follow up questions. It shows me that you're actually interested in the steps and the process and not just the solution. :)

OpenStudy (anonymous):

i am wanting to learn everything again i am trying to get into college i have been out of school for about 6 years now so its a little hard just jumping right back into this